r/maths•
SpazticSeal_69

[Probability] A question on drawing cards from a pack

I'm in my final year in secondary school in Ireland, we were given notes on probability lately. One example in the notes it is written that "The word 'and' means multiply, e.g pick 2 cards from 52, what is the probability of drawing a king and queen if you replace the drawn card. Ans. 4/52 * 4/52 "

I feel this is incorrect but I'm not sure. Surely the first draw there is 8/52 chance of drawing a king or queen, then the 2nd draw there's a 4/52 chance of drawing whatever card you didn't draw the first time.

8/52 * 4/52 = 2(4/52)^2

Or you can think about it that you can draw king then queen, (4/52 * 4/52) or queen then king (4/52 * 4/52) which is

2(4/52)^2

Am I misunderstanding the question or is my teacher oversimplifying?

Thanks for reading

Posted Dec 2, 2015, 10:36 PM UTC

136
Loading comments…